OpenStudy (anonymous):

The arch of a train tunnel is in the shape of a parabola. The maximum height of the tunnel is 20 m. If the origin of a grid is placed at one end of the arch where the tunnel contacts the road, the other point of contact with the road is 40 m away. Which quadratic function can be used to model the arch?

OpenStudy (whpalmer4):



OpenStudy (whpalmer4):

pfft! we don't need no steekin' answers to figure this out.

OpenStudy (whpalmer4):

A parabola is symmetrical about the vertex, so the high point of the arch (aka the vertex) is 1/2 of the way between 0 and 40. That means the vertex is at (20,20).

OpenStudy (campbell_st):

well the line of symmetry is x = 20 and the point (20, 20) is the vertex so the equation is y =a(x(x-40) or y = a(x^2 - 40x) sunstitute (20, 20) to find the value of a 20 = a(20^2 - 40(20)) 20 = -400a solve for a and then substitute into y = a(x^2 - 40x)

OpenStudy (anonymous):

thanks!